# Prime and Reach: Synthesising Body Motion for Gaze-Primed Object Reach

### v1

@@ 4 Method
4.1 Objective
Here, we address the task of goal-conditioned human motion generation with the ability to prime and reach a given object. Specifically, the task aims to generate human motion sequences $\{x^{i}\}_{i=1}^{N}$ of length $N$, where $x^{i}\in R^{J\times 3}$ represents 3D positions of $J$ body joints, guided either by desired goal pose or target goal object location as a condition. We consider and compare the two conditions. As sequences can be of varying length, we opt to predict a fixed-length sequence by uniformly up-/down-sampling $N$ frames. Predicted sequences can then be up- or down-sampled temporally so that the motion length is the same as the ground truth.
@@ 4.2 Prime & Reach Motion Diffusion Model
Conditioning. Diffusion models have demonstrated exceptional capability for text-conditioned motion generation [tevet2023mdm, zhang2023remodiffuse]. Motivated by this, we use a diffusion generative model, as in [tevet2023mdm] for our task. We present our architecture in Fig. 3. Starting from pure noise at $t=T$, the transformer decoder generates motion through iterative denoising over multiple diffusion timesteps $t=\{T,...,0\}$ where $t=0$ produces the predicted motion. This generation is guided through a set of conditions injected into the decoder. We condition our prime and reach motion generation on –
@@
  • Initial state of the body describing where and how the motion initiates. As P&R sequences do not start from a static or neutral pose, but are sampled from within a longer sequence, it is important to feed the initial pose and the velocity at the first frame, as this impacts the guided motion. We represent this by the starting pose $(\hat{x}^{1})\in R^{J\times 3}$ and the joint velocities i.e. $(\hat{x}^{1}-\hat{x}^{0})$ where $\hat{x}^{0}$ is the preceding frame before the curated P&Rsequence. We flatten and concatenate the start pose and velocities into a single vector, which forms our initial state.
@@
  • Goal. We evaluate two possible goal formulations. The first is the complete goal/target pose at the end of the motion $(\hat{x}^{N})$. The goal pose not only guides the motion to reach the object but additionally guides where the agent would stand relative to the object (through the full pose at the end of the motion) as well as which hand would be reaching the object (guided through the position of the hand joints). Second, we use the more challenging goal of only specifying the object location $(o_{3D}\in R^{3})$ as a condition. For this goal, the model needs to estimate where the body should be relative to the object and which hand to be the one reaching for the object.
@@
Architecture. Next, we explain how we use these conditions for the generation process. First, we encode the input text prompt $\mathbf{c}$ using a pre-trained text encoder [sanh2019distilbert]. Both the diffusion noise time step $t$ and $\mathbf{c}$ are then projected to a latent space and summed to get a token $\mathbf{z_{t}}$. Note that for generating motion conditioned on text, $\mathbf{z_{t}}$ is directly injected into the transformer decoder’s layers. We pre-train this model for text-conditioned motion generation. This pre-training enables the model to learn prior knowledge of fine-grained full-body motion involved in everyday activities.
@@
For learning P&R motion generation, we initialise our model with the pre-trained text-to-motion conditioned weights and fine-tune it with all three conditions. To add the initial state and goal condition, we first flatten and concatenate them into a single 1D vector. The resultant vector is linearly projected to a latent space to give $\mathbf{p}$. $\mathbf{z_{t}}$ is modified by adding condition $\mathbf{p}$ to $\mathbf{z_{t}}$.
$\mathbf{\tilde{z}_{t}}=\mathbf{z_{t}}+\mathbf{p}$ (3)
@@
This modified condition $\mathbf{\tilde{z}_{t}}$ is now injected into the transformer decoder’s layers through cross-attention blocks. This guides the motion sequence generation over multiple diffusion timesteps. We ablate other options of injecting p, including cross-attention, in the supplementary.
@@
Once de-noised, the decoder produces 263-dim representations of body joints $\{v^{n}\}_{n=1}^{N}$ where $v^{n}\in R^{263}$ combines local position, local rotation and local velocity of all 22 body joints. This is post-processed as $\{x^{n}\}_{n=1}^{N}=g(\{v^{n}\}_{n=1}^{N})$ to get the predicted 22 joint positions. We use the same $g$ as in [karunratanakul2023gmd, tevet2023mdm].
@@
Figure 3: P&R motion diffusion model for goal-conditioned motion generation. We concatenate the initial state of the human body and the goal pose/goal object as conditions, along with a text condition describing the type of action the motion is expected to perform. This accumulated condition is injected into the transformer decoder layers, which then outputs an NN-length motion sequence over multiple diffusion steps.
@@ 5 Experiments
Here we explain implementation details (Sec. 5.1), evaluation metrics (Sec. 5.2), baselines (Sec. 5.3), experimental results (Sec. 5.4) and ablations (Sec. 5.5).
@@ 5.1 Implementation Details
We pre-train our P&R motion diffusion model for text-conditioned motion generation on the large-scale Nymeria [ma2024nymeria] dataset to learn the motion prior of everyday activities. Nymeria provides large-scale full-body motion data of participants performing diverse actions captured by Xsens mocap sensors, accompanied by atomic narrations describing the actions. The narrations are used as text guidance. For pre-training, an initial learning rate of 1e41e-4 is used, for a maximum of 600K steps. We use a motion length of $N=150$ and classifier-free guidance with a probability of 0.2. The pre-training takes $\sim 36$ hours on one H200 GPU.
Initialised with the pre-trained weights, we fine-tune our Prime and Reach model on the curated P&R sequences from each dataset separately. This is because various datasets capture different types of activities. For fine-tuning, we use a learning rate of 5e55e-5 for 250K steps, taking around 15-20 hours on HD-EPIC and HOT3D. For MoGaze, ADT and GIMO, we use a learning rate of 1e51e-5 for $\sim 300K$, $\sim 150K$, and $\sim 170K$ steps respectively. We use $T=50$ diffusion steps in pre-training, finetuning, and inference, following [tevet2023mdm].
@@ 5.2 Evaluation Metrics
[…]
$\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\{\min_{t=t_{p}-\sigma}^{t_{p}+\sigma}(\cos^{-1}(\mathbf{\hat{H}_{t_{p}}},\mathbf{H_{t}}))\leq\theta\},$ (4)
where $B$ is the total number of motion sequences, $\mathbb{I}$ is the indicator function and $\sigma$, $\theta$ are hyperparameters. Unless mentioned otherwise, we use $\theta=16\deg$ and $\sigma=0.2$ sec. Prime Success gives the percentage of sequences where this angular error is $\leq$ threshold $\theta$. Having $\sigma$ relaxes the metric to allow predicted motion to prime the object in a temporal window of $2\sigma$ around $t_{p}$.
(2) Reach Success. We follow [diomataris2024wandr] and evaluate our reaching ability. This calculates the percentage of sequences where either wrist in the predicted motion reaches within 10 cm of the goal at $N$.
(3) Location Error. Also used in [karunratanakul2023gmd], we calculate the distance between the final pelvis locations of the ground truth and predicted motions and report the percentage of sequences where the distance is $\geq$ (50 cm).
@@
(6) Foot Skating, a common evaluation of the generated human motion [karunratanakul2023gmd], which measures the proportion of frames where either foot slides with velocity (>50>50 cm/s) between consecutive frames while maintaining contact with the ground (foot height <5<5 cm).
Table 2: Comparison of motion generation baselines on our curated P&R sequences using different metrics. We show results for HD-EPIC, MoGaze, HOT3D, ADT and GIMO separately. The baselines are grouped based on the kind of condition they use for generation. $\dagger$ denotes that the model is trained on HumanML3D. denotes the model is trained on Nymeria.[ma2024nymeria].
@@
No condition Static 30.15 15.02 0.82 50.92 0.45 1.64 4.17 1.03 74.26 0.64
Text MDM $\dagger$ 4.33 0.50 0.94 71.52 0.66 0.21 10.19 3.20 2.03 93.01 1.45 0.29
MDM 5.82 1.66 0.83 58.82 0.58 0.03 1.86 1.79 1.18 78.72 0.74 0.07
MDM 30.52 18.76 0.84 58.73 0.54 0.04 14.43 7.74 1.24 85.42 0.72 0.27
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 29.23 26.30 0.30 4.00 0.32 0.17 2.01 50.24 0.35 2.23 0.54 0.11
P&R 66.31 91.23 0.09 0.20 0.17 0.13 35.34 82.96 0.09 0 0.24 0.21
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 10.21 95.66 0.56 54.85 0.58 0.21 11.26 75.62 0.60 64.91 0.56 0.30
P&R 59.06 89.48 0.30 16.28 0.24 0.12 48.51 92.78 0.39 22.32 0.33 0.25
@@
No condition Static 42.13 31.38 0.30 6.96 0.27 7.29 11.46 1.98 72.92 1.12
Text MDM $\dagger$ 19.10 2.45 0.49 38.01 0.44 0.10 14.58 3.65 2.72 94.79 1.85 0.24
MDM 0.89 0.35 0.54 42.34 0.46 0.01 7.81 6.25 2.15 89.06 1.25 0.07
MDM 36.44 30.32 0.36 14.45 0.32 0.00 8.85 8.85 2.23 90.63 1.23 0.17
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 34.04 11.70 0.41 28.73 0.36 0.02 20.69 55.17 0.42 15.79 0.48 0.14
P&R 75.26 90.69 0.12 0.53 0.11 0 28.65 67.71 0.22 7.81 0.31 0.18
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 12.26 90.41 0.58 59.70 0.52 0.07 14.94 85.71 0.73 79.87 0.66 0.32
P&R 67.50 94.50 0.16 2.35 0.12 0 25.00 75.52 0.56 53.15 0.47 0.22
@@
No condition Static 14.28 0 2.85 100 1.84
Text MDM $\dagger$ 0 0 5.26 100 2.23 0.17
MDM 4.76 0 3.95 100 2.10 0.04
MDM 0 0 2.96 100 1.64 0.14
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 0 4.76 0.44 11.90 0.62 0.43
P&R 42.12 38.12 0.41 15.32 0.54 0.15
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 19.04 71.42 0.75 85.71 0.77 0.21
P&R 47.61 71.42 0.54 52.38 0.62 0.16

### v2

@@ 4 Method
Here, we address the task of goal-conditioned human motion generation with the ability to prime and reach a given object. Specifically, the task aims to generate human motion sequences $\{x^{i}\}_{i=1}^{N}$ of length $N$, where $x^{i}\in R^{J\times 3}$ represents 3D positions of $J$ body joints, guided either by desired goal pose or target goal object location as a condition. We consider and compare the two conditions.
@@ 4.1 Prime & Reach Motion Diffusion Model
Conditioning. Diffusion models have demonstrated exceptional capability for text-conditioned motion generation [84, 98]. Motivated by this, we use a diffusion generative model, as in [84] for our task. We present our architecture in Fig.˜3. Starting from pure noise at $t=T$, the transformer decoder generates motion through iterative denoising over multiple diffusion timesteps $t=\{T,...,0\}$ where $t=0$ produces the predicted motion. This generation is guided through a set of conditions injected into the decoder:
@@
  • Initial state of the body describes where and how the motion initiates. As our P&R sequences do not start from a static or neutral pose, but are sampled from within a longer sequence, it is important to feed the initial pose and the velocity, as this impacts the guided motion. We represent this by the starting pose $(\hat{x}^{1})\in R^{J\times 3}$ and the joint velocities i.e. $(\hat{x}^{1}-\hat{x}^{0})$ where $\hat{x}^{0}$ is the preceding frame before the curated P&R sequence. We concatenate the start pose and joint velocities as a single vector, which forms our initial state.
@@
  • Goal. In addition to the initial pose, motion generation expects the goal of the motion to be specified. We evaluate two possible goal formulations for P&R. The first is the final goal/target pose at the end of the motion $(\hat{x}^{N})$. The goal pose not only guides the motion to reach the object but additionally guides where the agent would stand relative to the object (through the full pose at the end of the motion) as well as which hand would be reaching the object (guided through the position of the hand joints). Second, we use the more challenging goal of only specifying the object location $(o_{\text{3D}}\in R^{3})$ as a condition. Given only the object location as goal, the model has more freedom to generate motions choosing the relative body pose to the object.
Importantly, we do not believe priming to be a condition. It should be implicitly learnt from the data. When using the goal pose, the direction where the object is, and thus how it can be primed is implicit. When using the goal object location, priming is synthesised by the body attending to this exact location.
Figure 3: P&R motion diffusion model for goal-conditioned motion generation. We concatenate the initial state of the human body and the goal pose/goal object as conditions, along with a text condition describing the type of action the motion is expected to perform. This accumulated condition is injected into the transformer decoder layers, which then outputs an NN-length motion sequence over multiple diffusion steps. At inference we perform diffusion latent noise optimisation over $M$ iterations using the same conditioning (i.e. initial state and goal pose or location).
@@
P&R motion generation with condition. Next, we explain how we use these conditions for the generation process. First, we encode the input text prompt $\mathbf{c}$ using a pre-trained text encoder [74]. Both the diffusion noise time step $t$ and $\mathbf{c}$ are then projected to a latent space and summed to get a token $\mathbf{z_{t}}$. Note that for generating motion conditioned on text, $\mathbf{z_{t}}$ is directly injected into the transformer decoder’s layers. We pre-train this model for text-conditioned motion generation. This pre-training enables the model to learn prior knowledge of fine-grained full-body motion involved in everyday activities.
@@
For learning P&R motion generation, we initialise our model with the pre-trained text-to-motion model weights and fine-tune it with all three conditions. To add the initial state and goal condition, we first flatten and concatenate them into a single 1D vector. The resultant vector is linearly projected to a latent space to give $\mathbf{p}$. $\mathbf{p}$ is injected through residual addition to the text condition $\mathbf{z_{t}}$ as $\mathbf{\tilde{z}_{t}}=\mathbf{z_{t}}+\mathbf{p}$.
@@
This allows the additional conditions to modulate the global context without introducing new attention pathways (e.g., cross-attention) that could disrupt the pre-trained text-conditioned motion prior (ablated in supplementary). This modified condition $\mathbf{\tilde{z}_{t}}$ is injected into the transformer decoder’s layers through cross-attention blocks and guides the generation over multiple diffusion timesteps.
@@
Once de-noised, the decoder produces 263-dim representations of body joints $\{v^{n}\}_{n=1}^{N}$ where $v^{n}\in R^{263}$ combines local position, local rotation, and local velocity of all 22 body joints. This is post-processed as $\{x^{n}\}_{n=1}^{N}=g(\{v^{n}\}_{n=1}^{N})$ to get the predicted 22 joint positions. We use the same joints-to-latent dimension conversion function $g$ as in [39, 84].
4.2 Training and Inference of P&R Model
@@
[…]
We then calculate the optimisation loss, $\mathcal{L}_{\text{opt}}=\mathcal{L}_{\text{init}}+\mathcal{L}_{\text{goal}}+\alpha\mathcal{L}_{\text{jerk}}$ where $\mathcal{L}_{\text{jerk}}$ controls the quality of the motion following [101]. This optimisation is iteratively performed over $M$ iterations (right figure in Fig.˜3).
@@ 5 Experiments
Here we explain implementation details (Sec.˜5.1), evaluation metrics (Sec.˜5.2), baselines (Sec.˜5.3), experimental results (Sec.˜5.4) and ablations (Sec.˜5.5).
@@ 5.1 Implementation Details
We pre-train our P&R motion diffusion model for text-conditioned motion generation on the large-scale Nymeria [57] dataset to learn the motion prior of everyday activities. Nymeria provides large-scale full-body motion data of participants performing diverse actions, including some priming and reaching activities, captured by Xsens mocap sensors, accompanied by atomic narrations describing the actions. This makes it a better pre-training dataset for P&R compared to the alternative HumanML3D [24]. The narrations are used as text guidance. For pre-training, an initial learning rate of 1e41e-4 is used, for a maximum of 600K steps. We use a motion length of $N=150$ and classifier-free guidance with a probability of 0.2. The pre-training takes $\sim 36$ hours on one H200 GPU.
Initialised with the pre-trained weights, we fine-tune our P&R model on the training split of our curated P&R dataset, training a single model on 16.7K P&R sequences. For fine-tuning, we use a learning rate of 5e55e-5 for 250K steps, which takes approximately 15-20 hours. We use $T=50$ diffusion steps in pre-training, fine-tuning, and inference, following [84]. During inference, we optimise the latent noise over $M=400$ iterations with an initial learning rate of 0.05 during $\mathcal{L}_{\text{opt}}.$ Following [16], we use $\alpha=0.5$.
@@ 5.2 Evaluation Metrics
(1) Prime Success. This is evaluating whether the generated motion is priming the target location before reaching it. Specifically, within a generous temporal window around the ground truth priming time $t_{p}$, we evaluate if the predicted motion exhibits intentional priming behaviour. While the generated motion can prime the object at a different time than the ground-truth, priming should happen well before reaching, not too far from the priming time. The intentional priming behaviour requires moving one’s gaze to look at the object, then attending to it. We formulate this as follows: 1) the head forward vector of the prediction motion ($\mathbf{\hat{H}}$) intersects the target location $o_{\text{3D}}$ within a proximity threshold for a minimum duration, and 2) the angular velocity of the head forward vector decelerates into the target location, signifying intentional priming from elsewhere. The reach metric per sequence is calculated as follows:
$\mathbb{I}\{\exists t\in[t_{p}-\sigma,t_{p}+\sigma]\mid\big(\forall k\in[t,t+\tau],dist(\mathbf{\hat{H}}_{k},o_{\text{3D}})\leq\delta\big)\land\Delta{\omega}_{t}<0\},$ (3)
where $\mathbb{I}$ is the indicator function and $dist(:,:)$ computes the orthogonal distance between $o_{\text{3D}}$ and the head forward vector. The term $\Delta{\omega}_{t}$ represents the change in angular velocity of the head forward vector at the onset of intersection; a negative value ($\Delta{\omega}_{t}<0$) indicates the active deceleration toward the target. We use $\tau=0.1$ sec, $\delta=25$ cm, and $\sigma=1.0$ sec. Having $\sigma$ relaxes the metric to allow predicted motion to prime the object in a temporal window of $2\sigma$ around $t_{p}$. The prime success then considers the percentage of sequences where an object is deemed primed.
(2) Reach Success. Following [15], this is the percentage of predicted sequences where either wrist reach within 10 cm of the goal.
(3) Location Error. Following [39], this is the percentage of motions, where final pelvis location of prediction is $\geq$ 50 cm away from ground truth pelvis.
@@
(6) Foot Skating, a common evaluation of the generated human motion [39], which measures the proportion of frames where foot slides while maintaining contact with the ground.
Table 2: Comparison of motion generation baselines on our curated P&R sequences using different metrics. While we train a single model for all datasets, we separate results per dataset. We show results for test splits of HD-EPIC, MoGaze, HOT3D, ADT, and GIMO separately. The baselines are grouped by the type of conditioning used for generation. $\dagger$ denotes the zero-shot inference. For MDM, we evaluate two pre-trained models: (1) trained on HumanML3D [24], and (2) trained on Nymeria.[57], denoted as $\ddagger$ and , respectively. Entries without a marker correspond to models fine-tuned on our P&R sequences.
@@
No condition Static 0.00 23.16 0.70 50.91 0.45 0.00 2.56 1.06 75.99 0.62
Text MDM $\dagger\ddagger$ 9.53 13.94 1.14 81.75 0.96 0.16 2.85 2.20 2.03 94.11 1.45 0.39
MDM $\dagger^{*}$ 9.52 13.47 0.85 62.13 0.59 0.06 3.11 1.85 1.19 79.76 0.73 0.06
MDM 12.40 18.76 0.84 58.73 0.54 0.04 5.27 2.90 1.24 85.42 0.72 0.05
+ Initial State & Goal Pose GMD $\dagger$[39] 33.27 30.77 0.26 2.00 0.32 0.10 8.11 10.86 0.35 2.23 0.54 0.11
GMD[39] 39.47 32.36 0.27 5.20 0.31 0.06 22.11 21.66 0.30 1.15 0.55 0.29
DNO $\dagger$[38] 43.42 87.44 0.07 2.60 0.30 0.04 17.39 69.71 0.09 6.81 0.64 0.07
DNO [38] 48.30 87.44 0.05 0.51 0.27 0.07 29.27 34.21 0.09 0.55 0.62 0.16
DartControl $\dagger$ [101] 30.06 81.23 0.14 2.20 0.43 0.15 38.26 71.16 0.11 0.14 0.52 0.35
DartControl [101] 35.29 88.27 0.11 1.86 0.38 0.09 35.07 62.61 0.12 0.14 0.60 0.51
P&R 52.75 97.40 0.09 1.58 0.21 0.03 45.24 96.47 0.05 0.00 0.32 0.16
+ Initial State & Object Loc. WANDR $\dagger$[15] 33.28 80.92 0.47 40.35 0.42 0.11 31.74 96.81 0.62 60.14 0.61 0.25
WANDR [15] 31.92 75.16 0.54 47.65 0.50 0.16 40.87 98.26 0.68 63.77 0.64 0.25
DNO $\dagger$[38] 37.34 100.00 0.47 37.87 0.44 0.05 27.24 100.00 0.64 59.56 0.87 0.09
DNO [38] 45.42 100.00 0.41 34.64 0.27 0.06 31.69 100.00 0.74 79.13 0.96 0.29
DartControl $\dagger$ [101] 28.82 89.20 0.47 40.82 0.52 0.13 42.90 100.00 0.54 50.29 0.72 0.40
DartControl [101] 30.04 89.42 0.44 34.16 0.47 0.08 40.14 100.00 0.56 54.35 0.74 0.54
P&R 51.00 100.00 0.38 25.26 0.27 0.03 62.37 100.00 0.46 36.43 0.56 0.11
@@
No condition Static 0.00 26.43 0.35 22.01 0.32 0.00 9.90 1.86 69.27 1.03
Text MDM $\dagger\ddagger$ 27.28 6.25 1.11 75.85 0.89 0.31 9.38 5.21 2.54 97.40 1.76 0.35
MDM $\dagger^{*}$ 8.65 4.30 0.51 36.20 0.44 0.00 3.65 6.25 1.98 83.33 1.16 0.06
MDM 19.25 30.32 0.36 14.45 0.32 0.00 10.58 15.29 2.23 90.63 1.23 0.17
+ Initial State & Goal Pose GMD $\dagger$[39] 31.85 20.16 0.38 25.00 0.37 0.02 16.47 12.94 0.35 10.58 0.48 0.21
GMD [39] 44.08 33.19 0.37 8.87 0.38 0.03 21.18 23.53 0.37 9.41 0.41 0.24
DNO $\dagger$[38] 45.83 90.18 0.05 9.67 0.23 0.02 25.88 52.94 0.04 5.88 0.56 0.08
DNO [38] 54.30 93.41 0.04 1.15 0.19 0.02 28.23 80.00 0.04 1.00 0.45 0.16
DartControl $\dagger$ [101] 40.51 89.77 0.17 2.15 0.24 0.01 14.12 67.06 0.22 8.24 0.62 0.24
DartControl [101] 45.76 89.77 0.18 2.02 0.28 0.02 20.00 56.47 0.25 10.59 0.66 0.29
P&R 58.65 99.19 0.04 0.00 0.14 0.00 35.42 98.82 0.02 0.00 0.37 0.12
+ Initial State & Object Loc. WANDR $\dagger$[15] 34.32 91.66 0.33 17.77 0.27 0.05 7.65 82.94 0.66 61.76 0.60 0.22
WANDR [15] 45.09 83.58 0.43 31.49 0.31 0.06 30.59 80.00 0.79 70.59 0.64 0.28
DNO $\dagger$ [38] 46.77 100.00 0.41 40.86 0.36 0.04 23.53 100.00 0.64 61.18 0.70 0.10
DNO [38] 63.04 100.00 0.35 22.17 0.31 0.05 37.64 100.00 0.62 60.78 0.61 0.10
DartControl $\dagger$ [101] 46.43 97.58 0.38 19.78 0.34 0.01 25.88 97.65 0.69 64.71 0.86 0.29
DartControl [101] 48.59 97.71 0.32 8.88 0.30 0.00 20.00 96.47 0.68 62.35 0.79 0.32
P&R 68.32 100.00 0.40 11.90 0.29 0.01 52.08 100.00 0.57 52.54 0.61 0.11
@@
No condition Static 0.00 0.00 3.41 100.00 1.86
Text MDM $\dagger\ddagger$ 0.00 0.00 3.69 100.00 2.17 0.45
MDM $\dagger^{*}$ 0.00 0.00 3.34 100.00 1.82 0.08
MDM 0.00 0.00 2.96 100.00 1.64 0.14
+ Initial State & Goal Pose GMD $\dagger$[39] 13.63 4.76 0.44 11.90 0.62 0.43
GMD [39] 27.27 9.09 0.45 4.54 0.61 0.25
DNO $\dagger$ [38] 18.18 27.27 0.15 9.09 0.80 0.18
DNO [38] 31.82 72.72 0.09 4.54 0.65 0.15
DartControl $\dagger$ [101] 9.52 14.29 0.33 9.52 1.37 0.07
DartControl [101] 19.05 14.29 0.32 9.52 0.84 0.07
P&R 50.00 90.90 0.03 0.00 0.51 0.10
+ Initial State & Object Loc. WANDR $\dagger$[15] 14.29 61.90 0.51 57.14 0.79 0.44
WANDR [15] 9.52 52.38 0.65 71.43 0.77 0.42
DNO $\dagger$ [38] 27.27 100.00 0.76 63.63 0.98 0.11
DNO [38] 40.91 100.00 0.39 40.91 0.72 0.11
DartControl $\dagger$ [101] 4.76 76.19 0.49 61.90 1.51 0.10
DartControl [101] 19.05 76.19 0.55 80.95 0.98 0.15
P&R 40.91 100.00 0.46 40.91 0.65 0.25